Zable Corporation has two divisions: Town Division and Country Division. The following report is for the most recent operating period.

Town Division Country Division

Sales $235,000 $192,000

Variable Expenses $103,400 $105,600

Traceable Fixed Expenses $80,000 $58,000

The company's common fixed expenses total $68,320

Required:

a. Prepare a contribution format income statement?

b. What is the Town Division break-even in sales dollars? Show all work!

c. What us the Country Division break-even in sales dollars? Show all work!

d. What is the company's overall break-even in sales dollars? Show all work!
Answer #1

A. Income statement


Total Company

Town Division

Country Division

Sales

$

427,000

$

235,000

$

192,000

Variable expenses

209,000

103,400

105,600

Contribution margin

218,000

131,600

86,400

Traceable fixed expenses

138,000

80,000

58,000

Segment margin

80,000

$

51,600

$

28,400

Common fixed expenses

68,320

Net operating income

$

11,680

B.

Town Division break-even:

Segment CM ratio = Segment contribution margin / Segment sales

= $131,600 / $235,000 = 0.560

Dollar sales for a segment to break even = Traceable fixed expenses / Segment CM ratio

= $80,000 / 0.560 = $142,857

C.

Country Division break-even:

Segment CM ratio = Segment contribution margin 梅 Segment sales
= $86,400 / $192,000 = 0.450

Dollar sales for a segment to break even = Traceable fixed expenses 梅 Segment CM ratio
= $58,000 / 0.450 = $128,889

D.

The company's overall break-even sales:

CM ratio = Contribution margin / Sales

= $218,000 / $427,000 = 0.51%

Total fixed expenses = Total traceable fixed expenses + Common fixed expenses

= $138,000 + $68,320 = $206,320

Dollar sales to break even = Total fixed expenses / CM ratio

= $206,320 / 0.51 = $404,549